# Finite size effects and optimization of the calculation of the surface tension in surfactant mixtures at liquid/vapour interfaces

### v1

@@ 2 Computational model
Different surfactant molecules were simulated, the sodium dodecyl sulfate (SDS), the dodecyltrimethylammonium bromide (DTAB), the cetyltrimethylammonium bromide (CTAB) and the zwitterionic lauryl sulfobetaine (SB-3), called Betaine in figure 1. The first one is anionic, the next two are cationic and the last one is nonionic surfactant, respectively (see figure 1). In all cases the atoms of the surfactant headgroups were explicitly simulated whereas the CH groups in the hydrocarbon tails were modelled using the united atom model (see figure 1). Counterions, Na and Br, were also introduced into the water phase to neutralize the SDS and DTAB (and CTAB) systems. For water, the TIP4P/ model was used [22] and all the unlike interactions were handled with the Lorentz-Berthelot rules [23].
@@ 3.2 Monolayers with surfactant mixtures
With the current surfactant LJ parameters, three monolayers were prepared with mixtures of CTAB/SDS, CTAB/DTAB and CTAB/Betaine at distinct compositions, 75/25, 50/50 and 25/75. Initial simulations were conducted with the same box dimensions as described in the previous section, i.e., nm and nm using an area per molecule of 26 Å/molecule for the CTAB/SDS and 40 Å/molecule for the CTAB/DTAB and CTAB/Betaine mixtures. The area per molecule is defined as the interfacial area divided by the number of surfactants ( / Num. surfactants). With these simulation boxes the surface tension was calculated, although at these conditions the results lead to large errors. Even negative values were found for the asymmetric compositions, which have no physical meaning (figure 2a). Exploring possible sources of errors, from the snapshots of the last configuration of the monolayers at small interfacial areas it is observed that they are distorted, i.e., they are not planar (top of figure 3). Therefore, equation (3.1) cannot be used, or in other words its application yields the not-physical results. Interesting are the surfactant molecules in the bulk water phase forming micelle-like structures for some surfactant compositions (bottom of figure 3).
@@ 3.3 Density profiles for the monolayers of surfactant mixtures
The results for the mass density profiles are shown in figures 5–7. The headgroup profiles [OSO for SDS and N(CH) for CTAB] are narrower for the systems with large box lengths than those with small lengths indicating that the monolayers at large interfacial areas are nearly planar. The monolayers in small areas, are not planar and then equation (3.1) cannot be used or incorrect values for the surface tension could be obtained.
@@ 3.4 Order of the surfactant tails and tilt angle
Additional information about the structure of the monolayers with small and large box areas can be obtained from an order parameter which tells us how the surfactants are arranged within them. In experiments, the ordering of the surfactant hydrocarbon tails is characterized by the so-called deuterium order parameter, , related to the average inclination of the C–H bond, in the CH groups, with respect to the surface normal. However, in computer simulations the order parameter is calculated, in the united atom model, by the following equation [30],

### v2

@@ 2 Computational model
Different surfactant molecules were simulated, the sodium dodecyl sulfate (SDS), the dodecyltrimethylammonium bromide (DTAB), the cetyltrimethylammonium bromide (CTAB) and the zwitterionic lauryl sulfobetaine (SB-3), called Betaine in figure 1. The first one is anionic, the next two are cationic and the last one is nonionic surfactant, respectively (see figure 1). In all cases the atoms of the surfactant headgroups were explicitly simulated whereas the CHn groups in the hydrocarbon tails were modelled using the united atom model (see figure 1). Counterions, Na+ and Br-, were also introduced into the water phase to neutralize the SDS and DTAB (and CTAB) systems. For water, the TIP4P/ model was used [22] and all the unlike interactions were handled with the Lorentz-Berthelot rules [23].
@@ 3.2 Monolayers with surfactant mixtures
With the current surfactant LJ parameters, three monolayers were prepared with mixtures of CTAB/SDS, CTAB/DTAB and CTAB/Betaine at distinct compositions, 75/25, 50/50 and 25/75. Initial simulations were conducted with the same box dimensions as described in the previous section, i.e., nm and nm using an area per molecule of 26 Å2/molecule for the CTAB/SDS and 40 Å2/molecule for the CTAB/DTAB and CTAB/Betaine mixtures. The area per molecule is defined as the interfacial area divided by the number of surfactants ( / Num. surfactants). With these simulation boxes the surface tension was calculated, although at these conditions the results lead to large errors. Even negative values were found for the asymmetric compositions, which have no physical meaning (figure 2a). Exploring possible sources of errors, from the snapshots of the last configuration of the monolayers at small interfacial areas it is observed that they are distorted, i.e., they are not planar (top of figure 3). Therefore, equation (3.1) cannot be used, or in other words its application yields the not-physical results. Interesting are the surfactant molecules in the bulk water phase forming micelle-like structures for some surfactant compositions (bottom of figure 3).
@@ 3.3 Density profiles for the monolayers of surfactant mixtures
The results for the mass density profiles are shown in figures 5–7. The headgroup profiles [OSO for SDS and N(CH3) for CTAB] are narrower for the systems with large box lengths than those with small lengths indicating that the monolayers at large interfacial areas are nearly planar. The monolayers in small areas, are not planar and then equation (3.1) cannot be used or incorrect values for the surface tension could be obtained.
@@ 3.4 Order of the surfactant tails and tilt angle
Additional information about the structure of the monolayers with small and large box areas can be obtained from an order parameter which tells us how the surfactants are arranged within them. In experiments, the ordering of the surfactant hydrocarbon tails is characterized by the so-called deuterium order parameter, , related to the average inclination of the C–H bond, in the CHn groups, with respect to the surface normal. However, in computer simulations the order parameter is calculated, in the united atom model, by the following equation [30],